OpenStudy (anonymous):

( 2 points) How much potassium chloride is produced, in grams, from a reaction that started with of 3.00 g of KClO3? (Assume all the KClO3 decomposes, and make sure the equation is balanced!!!) KClO3 → KCl + O2

OpenStudy (anonymous):

thanks for helping again!

OpenStudy (aaronq):

First balance the equation. Then convert the mass given to moles. -------------------------------------------------------- To interconvert between mass and moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. -------------------------------------------------- Next, use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\)

OpenStudy (anonymous):

uh give me a min trying to understand haha

OpenStudy (aaronq):

sure. all these stoichiometry problems are the same, so once you understand this one you can do all of them.

OpenStudy (anonymous):

i know the equation but my professor sucks at explaining everything and he uses other university power point and he gets lost and all that stuff so i am just trying to learn

OpenStudy (aaronq):

ok cool. we'll if you have doubts after you've tried it just gimme a shout

OpenStudy (anonymous):

2 2 3 for the balancing

OpenStudy (anonymous):

0.024 moles of KCLO3

OpenStudy (anonymous):

now what do i do?

OpenStudy (aaronq):

now you use a ratio to relate them. \(\dfrac{0.024}{2}=\dfrac{n_{KCl}}{2}\) you would solve algebraically, but since they coefficients are the same, then theres no work to be done (the coefficients cancel out). so now that you have the moles of KCl convert to grams

OpenStudy (anonymous):

so just divide 0.024/2 ?

OpenStudy (anonymous):

then multiply bby 2?

OpenStudy (aaronq):

that would be to find the moles of KCl, yes. But since you're multiplying and dividing by the same number, its redundant.

OpenStudy (anonymous):

ops i am wrong i would end up with the moles

OpenStudy (anonymous):

can you please explain ? sorry

OpenStudy (aaronq):

sure, the process in these stoichiometry problems is to convert whatever you're given to moles, then use the coefficients (via the ratio) to find the moles of what you're looking for. THEN you convert the moles to whatever they're asking for (grams, volume, etc.) So you have \(\dfrac{0.024}{2}=\dfrac{n{KCl}}{2}\rightarrow n_{KCl}=\dfrac{2*0.024}{2}=0.024 ~moles~of ~KCl\) now you convert these to what they asked for.

OpenStudy (anonymous):

thats the moles for KCLO3 though wait can i just divide the O3 moles from KCL? and get the answer?

OpenStudy (anonymous):

its asking for potassium chloride

OpenStudy (anonymous):

the moles is wrong its 0.0176

OpenStudy (anonymous):

wait thats wrong too

OpenStudy (anonymous):

the right moles is 0.012

OpenStudy (aaronq):

can i just divide the O3 moles from KCL? and get the answer? no, you can't do that. Is 0.012 moles for KClO3?

OpenStudy (anonymous):

yes

OpenStudy (aaronq):

okay, so from the balanced equation we know that the coefficients are the same. So the moles of KClO3 are equal to the moles of KCl. (This is apparent when you look at the ratio above) So now that you know the moles of KCl, convert them to mass

OpenStudy (anonymous):

uhhh yes i guess because its 2 2 3

OpenStudy (anonymous):

i am so lost sorry!! i know this the moles part

OpenStudy (aaronq):

it's because it's 2:2. IN comparison, if you were asked to find the moles of \(O_2\) produced, then it'd be \(\dfrac{0.012}{2}=\dfrac{n_{O_2}}{3}\)

OpenStudy (anonymous):

right

OpenStudy (aaronq):

so i checked and 0.024 moles was the right amount of mole of KClO3

OpenStudy (anonymous):

can't be because after you balance and multiply 2*39+2*35.45*16*6

OpenStudy (anonymous):

then divide 3 by that

OpenStudy (anonymous):

35.45+16*6

OpenStudy (aaronq):

you dont include the coefficient into the formula it;s just \(KClO_3\), not \(2~times~KClO_3\)

OpenStudy (anonymous):

oh yeah i was right the first time then

OpenStudy (anonymous):

so (0.024/2)= (1/2)

OpenStudy (anonymous):

because moles of KCL is just 1 on the other side

OpenStudy (aaronq):

hm no you dont need to divide by 2. It;s just algebra, \(\dfrac{0.024}{2}=\dfrac{x}{2}\) solve for x \(x=0.024\)

OpenStudy (aaronq):

you actually dont know the moles of KCl, so it's not 1. it's an unknown

OpenStudy (anonymous):

so thats it?

OpenStudy (anonymous):

0.024 is the moles of KCL?

OpenStudy (anonymous):

and convert to grams now?

OpenStudy (aaronq):

thats the moles, yeah. (this is useful when the coefficients arent the same) now convert to mass

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

so 0.024*122.45

OpenStudy (anonymous):

i mean 0.024 * 74.45

OpenStudy (aaronq):

yeah, the bottom one is right

OpenStudy (anonymous):

so 1.79 g THANK YOU SO MUCH! I CONFUSED MYSEL

OpenStudy (anonymous):

can you help me with one more problem? this is the last one

OpenStudy (aaronq):

yeah, its confusing at first. you'll be fine after you do a few of them. and yeah, post it

OpenStudy (anonymous):

okay about to i will post a screenshot of this because it has graph